# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

### v1

@@ 4.3 Experiments
Figure 8: Complexity tuning results for optimal decision trees of (a) $\operatorname{max-depth}=4$ and (b) $\operatorname{max-depth}=5$ for five runs on (a) 164 data sets and (b) 131 data sets.
Figure 9: Nemenyi critical distance rank test for ODTs of maximum depth four. The average rank per tuning method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
In our experiments we aim to answer the following question: how do the ODT tuning methods compare in out-of-sample accuracy?
We tune each method using five-fold cross-validation on the 165 OpenML sets data described in Section 3.4.1.444The code will be made public on paper publication. Regardless of the tuning method, we impose a maximum depth constraint of four or five. For runtime comparison, we ran all experiments on an Intel Xeon E5-6248R 3.0GHz with 100GB RAM using one thread. All tuning methods were implemented in STreeD (van der Linden et al., 2023).
@@
Figs. 8 and 8 show the performance of the complexity tuning method on the OpenML data sets. We exclude data sets from the analysis if any method exceeded the one-hour time out: for depth four, one data set out of 165 is removed, and for depth five, 34 data sets. The results show that tuning the size performs best, although all methods are close.
Surprisingly, Fig. 9 shows that tuning only the depth is not statistically significantly worse even though it tests much fewer configurations than other approaches. On average it does yield larger trees since ODT methods will use as many nodes as allowed to maximize the objective on the training data. Tuning the minimum support also yields larger trees. In our experiments, tuning the question length works well for depth four but worse for depth five, but it is unclear why. Tuning through smoothing is also a new promising approach. Tuning the G-Test threshold and no tuning are significantly worse than tuning the size.
@@
Inspecting the results shows that the differences among tuning methods are largest for medium-sized data sets (several hundred to ten thousand samples). For example, on the 77 data sets with 250 or more and 10,000 or fewer samples, a Wilcoxon signed rank test shows tuning the size is statistically significantly better than tuning the depth (p-value $\approx 0.03$). From this, we hypothesize that for small data sets the validation sets are too small to yield good estimates of the test scores, whereas for large data sets, the risk of overfitting and thus the importance of tuning decreases.
Fig. 10 shows the runtime for a five-fold cross-validation over the parameter options described above for trees of maximum depth four. Note that these runtimes could easily be changed by considering more, less, or other possible parameter configurations, although this may also impact the accuracy performance. The left plot shows what percentage of computations for the five runs on the 165 data sets finished within the given runtime. The middle and right plots show the runtime for data sets for an increasing number of instances and binary features. Each marker represents the average runtime for one data set.
Apart from no tuning, which yields significantly worse accuracy results, tuning the depth or the size of the tree is the most efficient approach because of how STreeD, the ODT method used in the experiments, works. Since STreeD is a dynamic programming approach, it can reuse previously cached results when searching for increasingly large trees.
Figure 10: Runtime distributions for each tuning methods for $\operatorname{max-depth}=4$. The scatter plots show one marker for every data set averaged over five runs. Tuning the depth or the size of the tree is the most efficient tuning approach.
In conclusion, complexity tuning of ODTs is necessary. All previously used tuning approaches obtain similar accuracy results. Tuning the depth or the size of the tree (when using cached solutions) is more efficient than other approaches. Two new tuning approaches are promising: tuning the question length and smoothing. Future work could further explore combining multiple approaches, e.g., tuning the size and the smoothing level.
Based on our experiments in Sections 3 and 4, we recommend to train optimal decision trees as follows:
@@
    Optimizing the accuracy on average yields the best out-of-sample accuracy.
@@
    For small data sets, maximize a non-concave objective such as M-loss.
    For data sets with less than a few hundred samples, optimizing non-concave objectives such as M-loss performs better than maximizing accuracy as it has an added regularization effect.
@@
    Tuning the size, complexity cost, smoothing parameter, or depth, on average, yields similar out-of-sample results. However, tuning the depth yields larger trees and is worse for medium-sized data sets (a few hundred to ten thousand samples). Using DP approaches, tuning the size is more runtime efficient. For large data sets, tuning is less important, and tuning the depth is more runtime efficient.
5 Optimal Compared with Greedy
To understand the differences between greedy and optimal approaches for learning decision trees, we collected claims from the literature and evaluated these claims with extensive experiments. Below, we list the claims discussed in this section and summarize their verdicts:
@@ 5 Optimal Compared with Greedy
We evaluate the accuracy of depth three and four trees on 164 data sets and find an average improvement of 1.1% and 1.3% of optimal over greedy approaches.
@@
Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
@@
We evaluate the accuracy of trees with 1 to 16 nodes on 165 data sets and find that the size-weighted accuracy of optimal methods is, on average, higher than that of greedy methods. The size-weighted accuracy is a new metric we propose in Section 5.3 to measure the accuracy-interpretability trade-off.
@@
Experiments on synthetic and real data show that the performance of size-unconstrained greedy trees keeps improving with more samples, while also growing much larger trees than ODTs. In contrast, size-constrained greedy trees do not improve after some point and stay worse than optimal.
@@
With hyperparameter tuning, we do not find significant performance differences between optimal and greedy methods with small numbers of samples (up to 200) nor more sensitivity to noise. When trained with the same size constraint, optimal methods dominate with more samples.
@@
The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
@@
Our experiments on synthetic data with ground-truth trees of depths 1 to 5 show that the expected depth of all methods remains similar to the true depth.
@@ 5.1 Previous Comparisons
5.1.1 Comparisons in ODT Papers
5.1.2 Other Comparisons
Murthy and Salzberg (1995) compare the greedy approach to known true (optimal) trees on a synthetic benchmark. They observe that the greedy tree is approximately one standard deviation larger than the true tree size while the expected depth (question length) is very close to optimal, which was also observed by Goodman and Smyth (1988); Murphy and McCraw (1991) (Claim 6). However, the maximum depth of greedy trees is on average approximately two times higher than the true depth. When they increase the data set size linearly with the true tree complexity, they observe almost no drop in accuracy for the greedy approach (Claim 4). From this result, Costa and Pedreira (2023) hypothesize that the gap between optimal and greedy approaches diminishes for more data (Claim 3).
5.1.3 Discussion
@@
From the comparisons listed above, the following general observations can be made about how greedy and optimal methods are compared:
  • Both greedy and optimal methods are often not correctly tuned, or not tuned at all.
  • The papers that propose ODTs typically compare ODTs with greedy methods under the same size constraint, while other papers compare ODTs with a restricting size limit and other greedy methods without (or with a very permissive) a size limit.
  • Comparisons are often limited to a small number of small data sets.
  • Comparisons are sometimes limited to small trees.
We briefly discuss each of these in more detail.
When comparing with CART, the papers above show several modifications to how CART is trained and tuned. We assess the impact of each of these modifications on the accuracy in Section 5.2 below.
Several papers evaluated ODTs without tuning the size. As shown in Section 4.3, this significantly worsens the ODT performance. Others evaluate ODTs with a substantial restriction on the tree size, resulting in shallow underfitting trees.
Almost all the ODT papers compare with CART under the same tree-size constraint and draw a positive conclusion on the ODTs’ performance. The other papers typically compare ODTs under a size constraint with unconstrained non-optimal approaches and sometimes draw negative conclusions about the ODTs’ performance. Both comparisons are useful for different purposes. The central claim from the ODT papers is that ODTs have a better accuracy-interpretability trade-off. Therefore, we further discuss measuring this trade-off in Section 5.3 below.
@@ 5.2 Training CART
In the experiments reviewed above, ODTs are often compared to a modified version of CART, for example, to allow for a direct comparison under similar circumstances. We test the impact of these modifications to assess the validity of these previous comparisons and inform future comparisons. We assess the following typical modifications: 1) tuning the depth instead of the complexity cost; 2) binarizing the feature data; or 3) running CART while imposing an additional depth constraint.
We compare CART’s performance with these modifications against unmodified CART on the 165 OpenML data sets used before. We approximate the unconstrained CART with a maximum depth of 20. We set the constrained depth limit to four, because of its common use in ODT comparisons. As before, the binarized data has up to ten binary features per continuous feature by using thresholds on ten quantiles or one-hot encoding of categorical features with a maximum of ten categories.
We apply cost-complexity pruning as done in RPart (Therneau et al., 2023): we train a fully expanded tree and obtain the cost-complexity path with all possible cost-complexity values from that tree and use the geometric mean to get the midpoints of those values. We use five-fold cross-validation to find the best cost-complexity parameter among the midpoints and retrain a tree on the full training data with this parameter.
Fig. 11 shows CART’s performance under these modifications. The largest differences are between the depth-constrained and the unlimited depth variant. Binarization has only a small impact on the performance (this does not necessarily generalize for more coarse binarizations). When a strict depth limit is imposed, tuning the depth instead of the complexity cost has a small impact but for the unlimited depth case, this significantly hurts CART’s performance. Fig. 11 also shows significantly different tree sizes for CART’s modifications. Both binarization and depth tuning result in larger trees. From these results we can conclude the following best practices:
@@
    Training on binarized data does not significantly harm the performance, provided the binarization is not too coarse, but should be explicitly noted.
@@ 5.3 Accuracy-Interpretability Trade-Off
Figure 12: Typical accuracy-interpretability trade-off for greedy and optimal. Optimal has a slight advantage for small trees but both methods converge for large size limits.
ODT papers typically compare ODTs with greedy heuristics under a similar size constraint. This is partly motivated by the definition of ODTs because ODTs are defined as trees that maximize training performance under a given size limit. No theoretical claim is made about its out-of-sample performance or the performance without a size constraint. In fact, finding an optimal tree without a size constraint is trivial: it is obtained by splitting in any way until no further split can be made. It is precisely the size constraint that makes it NP-Hard (Hyafil and Rivest, 1976).
@@ 5.4 Experiments on Literature Claims
To evaluate the claims made in previous papers, we compare optimal and greedy methods on synthetic and real data sets. Our experiment setup follows the best practices introduced in Recommendations 2.555The code will be made public on paper publication.
@@ 5.4.1 Experiment Setup
We follow the synthetic data setup from Murthy and Salzberg (1995); Bertsimas and Dunn (2017) and Dunn (2018). We generate $n$ random training instances with $p$ numeric features, uniformly distributed over $[0,1]$. For a given percentage $f$ of the training data, we add feature noise by adding noise uniformly drawn from $[-0.25,0.25]^{p}$. We binarize the numeric features by threshold predicates on 10 quantiles per numeric feature. We generate a random binary tree on this binarized data of a maximum depth $d$ with at most $2^{d}$ leaf nodes. We choose random splits on the data such that each leaf node contains at least 5 instances. The binary labels of each leaf node are assigned alternately, such that no split leads to two leaf nodes with the same label. After this, we add class noise to a given percentage $c$ of the data by flipping its label. By default, we set the true tree depth to three, the number of instances $n=1000$, the number of numeric features $p=3$, the feature noise $f=0\%$, and the class noise $c=0\%$. For each training set, we create a corresponding test set without noise of $1000$ instances per leaf node in the generated tree.
[…]
We also evaluate the methods on the same OpenML data sets as in Section 3.4 and add two more datasets: covertype and Higgs. These additional datasets were chosen for their large number of samples, allowing us to investigate performance under various sample sizes.
We evaluate both CART and ODT on the binarized data to eliminate this difference between the two methods and focus only on the difference between greedy and optimal search. Comparing on binarized data is commonly done (Lin et al., 2020; Demirović et al., 2022). We show in Section 5.2 that for the data sets in our benchmark, the impact of this binarization on CART is only small. We train ODTs using STreeD by maximizing accuracy and tuning the size of the tree, according to our analysis in Section 3 and 4 (Recommendations 1). We train CART trees with and without a depth limit, using cost-complexity tuning, following the best practices mentioned in Section 5.2.
In our synthetic experiments, apart from the familiar test accuracy and number of leaf nodes, we also measure the following:
@@
The Question length (or expected depth) is the average number of branching nodes an instance visits when evaluated (lower is better).
Figure 13: Optimal (blue) versus CART (red) out-of-sample accuracy results on five runs for 164 data sets. CART without a depth limit performs similarly to optimal with a maximum depth of four. CART with the same depth limit performs significantly worse.
Figure 14: Nemenyi critical distance rank test for optimal versus CART. The average rank per method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
@@ 5.4.2 Out-of-Sample Accuracy
To evaluate Claim 1, we evaluate both the ODT approach and CART on the OpenML data sets with a depth limit of three and four. We also compare with CART without a depth limit. Since ODT with a maximum depth of four hits the time-out of one hour for one data set, we report the results for the other 164 data sets.
Fig. 13 and 14 show that ODT and greedy (without a depth limit) perform similarly on average, but the greedy approach yields much larger trees. When compared under the same depth constraint as stated in Claim 1, optimal outperforms greedy with an average improvement of 1.1% and 1.3% for depths three and four respectively. These differences are statistically significant. Since optimal algorithms optimize the complete decision tree, instead of greedily improving the tree, they can achieve better scores.
Fig. 15 shows the distribution of the differences between optimal and greedy when trained with maximum depth four. For small data sets ($n\leq 250$), the average advantage of the optimal approach is $0.7\%\pm 8.3$. Because data is sparse, both methods are sensitive to overfitting, resulting in some large accuracy differences, but on average, the optimal approach is better. For larger data sets ($n>250$), the average improvement of optimal over CART is $1.9\%\pm 4.7$. The difference is larger and the variance is lower. These results confirm Claim 1.

### v2

@@ 4.2 Experiments
The experiments aim to answer the following questions: how do the ODT tuning methods compare in out-of-sample accuracy and how do they respond to increasing noise?
@@
Figure 7: Complexity tuning results for ODTs of (a) $\operatorname{max-depth}=4$ and (b) $\operatorname{max-depth}=5$ for five runs on (a) 180 data sets and (b) 157 data sets.
4.2.1 Experiment setup
We evaluate each tuning method on the same 180 OpenML and synthetic data sets described in Section 3.4.1. Regardless of the tuning method, we impose a maximum depth constraint of four or five. For each tuning method, we select at most $k$ different parameter settings. For depth, these settings are selected using equal linear spacing between the $k$ options. For the others, the $k$ settings are selected using equal spacing in the log scale. In the results presented here, we set $k=16$ (because a depth-four tree can have zero to fifteen branching nodes, i.e., 16 options). We provide more details on how the values are chosen and experiments for other values of $k$ in Appendix B. All tuning methods are implemented in STreeD (van der Linden et al., 2023).
For most data sets, we obtain trees well within the maximum depth limit of five, and therefore we do not extend this analysis to larger depth limits.
4.2.2 Results on Real Data Sets
@@
Figs. 7 and 7 show the performance of the complexity tuning method on the OpenML data sets. We exclude data sets from the analysis if any method exceeded the two-hour time-out: for depth four, no data sets are removed, and for depth five, 23 data sets. Surprisingly, the results show that all tuning methods obtain similar accuracies: there is no statistically significant difference between any of the approaches. In terms of optimizing accuracy, the only conclusion is that using any of the tuning approaches is better than no tuning.
However, other differences can be observed between the methods. For example, tuning the question length, minimum support, or depth yields slightly larger trees. Furthermore, in Appendix B, we measure the runtime and test the performance of the methods for different numbers of parameter settings $k$. This shows that tuning only the depth of the tree is by far the fastest approach.
[…]
Figure 8: Relative test accuracy of ODT tuning methods compared to tuning the size (number of nodes) of the tree. No tuning, tuning the minimum support, and tuning the question length are significantly worse than the other tuning approaches when noise is present.
@@
Inspecting the results shows that the differences among tuning methods are largest for medium-sized data sets (several hundred to ten thousand samples). For example, on the 82 data sets with 250 or more and 10,000 or fewer samples, a Wilcoxon signed rank test shows tuning the size is statistically significantly better than tuning the depth (p-value $\approx 0.01$).
4.2.3 Results on Synthetic Data
To further test the tuning methods’ performance, we evaluate them on the synthetic data sets. Fig. 8 shows that as the amount of feature noise increases, tuning becomes increasingly more important. It also shows that tuning the question length is slightly worse than the other tuning methods. No clear differences can be observed between the other tuning methods.
Fig. 8 shows larger differences when increasing the amount of class noise. Again, not tuning is significantly worse when noise is present. Tuning the minimum support and the question length is significantly worse than the other tuning methods when class noise is present. At $50\%$ class noise, this difference obviously disappears because then all training labels are basically decided randomly.
Interestingly, Fig. 8 shows no significant difference in tuning only the depth versus the more expensive procedures of tuning the size, the cost-complexity, or the smoothing level.
4.3 Conclusion
In conclusion, complexity tuning of ODTs is necessary. On our real data sets, all previously used tuning approaches, obtain similar accuracy results. On the noisy synthetic data, on the other hand, tuning the question length or minimum support is significantly worse than the alternatives. Tuning the depth is more time efficient than other approaches, obtains the same accuracy, but does so with slightly more nodes. One new tuning approach is promising: tuning the amount of smoothing. Based on our experiments in Sections 3 and 4, we recommend training ODTs as follows:
@@ 4.3 Conclusion
    Optimizing the accuracy on average yields the best out-of-sample accuracy. Do not optimize concave proxies such as Gini impurity or entropy.
@@
    In noisy data sets, consider an objective with an additional regularizing effect.
    In noisy data sets, optimizing objectives such as the pessimistic binomial and minimum error may perform better than maximizing accuracy as they have an additional regularization effect by encouraging model sparsity.
@@
    Tuning the size, complexity cost, smoothing, or depth, on average, yields similar out-of-sample results. However, tuning the depth yields larger trees. For large data sets, tuning is less important, and tuning the depth is more runtime efficient.
5 Comparing Optimal and Greedy Decision Trees
To understand the differences between greedy and optimal approaches for learning decision trees, we collected claims from the literature and evaluated these claims with extensive experiments. Below, we list the claims discussed in this section and summarize our results:
@@ 5 Comparing Optimal and Greedy Decision Trees
We evaluate the accuracy of depth three and four trees on 180 data sets and find an average improvement of 1.3% and 1.0% of optimal over greedy approaches.
@@
Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
@@
We evaluate the accuracy of trees with 1 to 16 nodes on 180 data sets and find that the size-weighted accuracy of optimal methods is, on average, significantly higher than that of greedy methods. The size-weighted accuracy is a new metric we propose in Section 5.3 to measure the accuracy-interpretability trade-off.
@@
Experiments on synthetic and real data show that size-constrained greedy trees do not improve after some point and stay worse than optimal. In contrast, the performance of size-unconstrained greedy trees keeps improving with more samples, while also growing much larger trees than ODTs.
@@
With hyperparameter tuning, we do not find significant performance differences between optimal and greedy methods with small numbers of samples (up to 250) nor more sensitivity to noise.
@@
The question length of greedy trees remains (in practice) close to the optimal tree depth (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
@@
Our experiments on synthetic data show that the question length of optimal and greedy methods remains similar in (almost) all (practical) scenarios. Only for very noisy data, does CART yield much longer question lengths.
@@ 5.1 Previous Comparisons
In Appendix C, we list previous comparisons between greedy and optimal approaches. These comparisons can be grouped into ODT papers that propose new ODT methods, and other papers. From these comparisons, the following general observations can be made about how greedy and optimal methods are compared:
  • Both greedy and optimal methods are often not correctly tuned, or not tuned at all. When comparing with CART, many papers show several modifications to how CART is trained and tuned. We assess the impact of each of these modifications on the accuracy in Section 5.2 below. Additionally, several papers evaluated ODTs without tuning the complexity. As shown in Section 4.2, this significantly worsens the ODT performance. Others evaluate ODTs with a substantial restriction on the tree size, resulting in shallow underfitting trees.
  • Almost all ODT papers compare with CART under the same tree-size constraint and draw a positive conclusion on the ODTs’ performance. The other papers typically compare ODTs under a size constraint with unconstrained non-optimal approaches and sometimes draw negative conclusions about the ODTs’ performance. Both comparisons are useful for different purposes. One simply evaluates out-of-sample performance, while the other evaluates the claim that ODTs have a better accuracy-interpretability trade-off. We further discuss measuring this trade-off in Section 5.3 below.
@@ 5.2 Training CART
In the comparisons reviewed in Appendix C and summarized above, ODTs are often compared to a modified version of CART, for example, to allow for a direct comparison under similar circumstances. We test the impact of these modifications to assess the validity of these previous comparisons and inform future comparisons. We assess the following typical modifications: 1) tuning the depth instead of the complexity cost; 2) binarizing the feature data; or 3) running CART while imposing an additional depth constraint.
We compare CART’s performance with these modifications against unmodified CART on the 180 OpenML data sets used before. We approximate the unconstrained CART with a maximum depth of 20. We set the constrained depth limit to four, because of its common use in ODT comparisons. As before, the binarized data has up to ten binary features per continuous feature by using thresholds on ten quantiles or one-hot encoding of categorical features with a maximum of ten categories.
In addition to the depth limit, we apply cost-complexity pruning as done in RPart (Therneau et al., 2023): we train a fully expanded tree and obtain the cost-complexity path with all possible cost-complexity values from that tree and use the geometric mean to get the midpoints of those values. We use cross-validation to find the best cost-complexity parameter among the midpoints and retrain a tree on the full training data with this parameter.
Unlike RPart, we take the best performing cost-complexity parameter, and not the largest complexity cost which performs within one standard error of the best performing one. In our preliminary tests, this resulted in larger trees but better out-of-sample accuracy.
@@
Fig. 9 shows CART’s performance under these modifications. The largest differences are between the depth-constrained and the unlimited depth variant. Binarization has only a small impact on the performance (this does not necessarily generalize for more coarse binarizations). When a strict depth limit is imposed, tuning the depth instead of the complexity cost has a small impact but for the unlimited depth case, this significantly hurts CART’s performance. Fig. 9 also shows significantly different tree sizes for CART’s modifications. Both binarization and depth tuning result in larger trees. From these results, we can conclude the following best practices:
{mdframed}
Recommendations 3 (Training CART).
@@
    CART trained with a depth limit results in significantly different results than CART without a depth limit.
@@
    Tuning CART’s depth rather than the cost-complexity yields larger trees.
@@
    Training CART on binarized data should be clearly stated.
    Depending on the binarization, training on binarized data may or may not significantly harm the performance.
@@ 5.3 Accuracy-Interpretability Trade-Off
ODT papers typically compare ODTs with greedy heuristics under a similar size constraint. This is partly motivated by the definition of ODTs because ODTs are defined as trees that maximize training performance under a given size limit. No theoretical claim is made about its out-of-sample performance or the performance without a size constraint. In fact, finding an optimal tree without a size constraint is trivial: it is obtained by splitting in any way until no further split can be made.
Figure 10: Typical accuracy-interpretability trade-off for untuned greedy and optimal decision trees. ODTs have a slight advantage for small size limits but both methods converge for large size limits.
@@ 5.4 Experiments on Literature Claims
To evaluate the claims made in previous papers, we compare optimal and greedy methods on synthetic and real data sets, using the best practices introduced in Recommendations 1-3.
@@ 5.4.1 Experiment Setup
We again evaluate on both the OpenML and synthetic data sets introduced in Section 3.4.1. We add two more real data sets to evaluate Claim 3: covertype and Higgs. These two were chosen for their large number of samples, allowing us to investigate performance under various sample sizes (566,602 and 940,160 respectively). For the synthetic data, unless otherwise specified, we set the true tree depth to three, the number of instances $n=1000$, the number of numeric features $p=3$, the feature noise $f=0$, and the class noise $c=0\%$. We test with changing the number of instances ($n=50$, $100$, $250$, $1000$, $10000$), the number of features ($p=2$, $4$, $6$, $8$), and with changing the noise as done in Section 3.4.1. Additionally, we add synthetic data sets with a linear separator instead of a tree as the ground truth. The weights of the linear separator are chosen randomly from a normal distribution. We repeat each configuration 500 times and report averages over these 500 runs.
We evaluate CART and ODT on the binarized data to eliminate this difference between the two methods and focus only on the difference between greedy and optimal search. Comparing on binarized data is commonly done (Lin et al., 2020; Demirović et al., 2022). Section 5.2 shows that the impact of this binarization on CART is small for the data sets in our benchmark. In Appendix D.2, we also evaluate the effect of binarization for ODTs.
In our synthetic tree experiments, apart from the familiar test accuracy and number of leaf nodes, we also measure the following:
@@
The question length is the average number of branching nodes an instance visits when evaluated (lower is better).
@@ 5.4.2 Out-of-Sample Accuracy
Figure 11: Out-of-sample accuracy of CART and ODTs compared on five runs for 180 data sets. (a) Optimal (blue) versus CART (red). CART without a depth limit performs best but yields much larger trees. (b) Nemenyi critical distance rank test for optimal versus CART. The average rank per method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar. With the same depth limit, optimal performs significantly better than CART, confirming Claim 1.
@@
To evaluate Claim 1, we evaluate both the ODT approach and CART on the OpenML data sets with a depth limit of three and four. We also compare with CART without a depth limit. Fig. 11 shows that CART (without a depth limit) performs better (but not significantly) than the ODT approach with a depth-four depth limit, but CART yields much larger trees.
When compared under the same depth constraint, as stated in Claim 1, optimal significantly outperforms greedy with an average improvement of 1.3% and 1.0% for depths three and four respectively. Since optimal algorithms optimize the complete decision tree, instead of greedily improving the tree, they can achieve better scores.
Fig. 12 shows the distribution of the differences between optimal and greedy when trained with maximum depth four. For small data sets ($n\leq 250$), the average advantage of the optimal approach is $0.2\%\pm 0.4$ (mean $\pm$ standard error). Some large accuracy differences occur because the test sets for these small data sets are so small that a single misclassified instance can increase or decrease accuracy by several percent. For larger data sets ($n>250$), the average improvement of optimal over CART is $1.6\%\pm 0.2$. The difference is larger and the standard error is lower. These results confirm Claim 1.